OpenStudy (andriod09):

How do you solve: \[-\frac{1}{3}y-2=-6\]

OpenStudy (ajprincess):

Add 2 to both sides. then multiply by -3 both sides.

OpenStudy (ajprincess):

Does that help? @andriod09

OpenStudy (andriod09):

Sorry, had to get water. Kind of. I'm stuck at this: \[-\frac{4}{1}(-\frac{3}{1})\]

OpenStudy (ajprincess):

-4*-3=? (-)*(-)=(+)

OpenStudy (andriod09):

Oh, so you do multiply the tops and leave the bottoms alone?

OpenStudy (ajprincess):

\[\frac{4}{1}=4\]

OpenStudy (andriod09):

Don't I feel stupid. >.<

OpenStudy (ajprincess):

if a number is divided 1 then u get the same number as the answer

OpenStudy (ajprincess):

still stuck? Am sorry if i have confused u more.

OpenStudy (ajprincess):

@andriod09

OpenStudy (andriod09):

Sorry, browser crashed.

OpenStudy (ajprincess):

\[-\frac{1}{3}y-2=-6\] Add 2 both sides. then u will get \[-\frac{1}{3}y-2+2=-6+2\] \[-\frac{1}{3}y+0=-4\] \[-\frac{1}{3}y=-4\] Multiply by -3 both sides. \[-3*-\frac{1}{3}y=-4*-3\] \[-\cancel3*-\frac{1}{\cancel3}y=-4*-3\] \[-1*-\frac{1}{1}y=-4*-3\] \[-1*-1y=-4*-3\] \[1y=-4*-3\] \[y=-4*-3\] Does that help? @andriod09

OpenStudy (andriod09):

I already got it. Thanks.

OpenStudy (ajprincess):

Welcome:)